# PULASki: Learning inter-rater variability using statistical distances to improve probabilistic segmentation

### v1

@@ Abstract
In the domain of medical imaging, many supervised learning based methods for segmentation face several challenges such as high variability in annotations from multiple experts, paucity of labelled data and class imbalanced datasets. These issues may result in segmentations that lack the requisite precision for clinical analysis and can be misleadingly overconfident without associated uncertainty quantification. We propose the PULASki as a computationally efficient generative tool for biomedical image segmentation that accurately captures variability in expert annotations, even in small datasets. Our approach makes use of an improved loss function based on statistical distances in a conditional variational autoencoder structure (Probabilistic UNet), which improves learning of the conditional decoder compared to the standard cross-entropy particularly in class imbalanced problems. We analyse our method for two structurally different segmentation tasks ( intracranial vessel and multiple sclerosis (MS) lesion) and compare our results to four well-established baselines in terms of quantitative metrics and qualitative output. These experiments are characterised by class-imbalanced, low signal to noise ratio data sets with highly ambiguous features. Empirical results demonstrate the PULASKi method outperforms all baselines at the 5% significance level. Our experiments are also of the first to present a comparative study of the computationally feasible segmentation of complex geometries using 3D patches and the traditional use of 2D slices. The generated segmentations are shown to be much more anatomically plausible than in the 2D case, particularly for the vessel task. Our method can also be applied to a wide range of multi-label segmentation tasks and and is useful for downstream tasks such as hemodynamic modelling (computational fluid dynamics and data assimilation), clinical decision making, and treatment planning.
@@ 1 Introduction
The recent explosion in deep learning based medical image segmentation techniques has led to a dramatic improvement in the rapid diagnosis, treatment planning and modelling of various diseases li2021multiple ; ni2020global . Medical image segmentation has specific challenges compared to other classical computer vision tasks, which has spurred the development of several bespoke machine learning tools chatterjee2022strega ; antonelli2022medical ; qureshi2023medical . Such issues include 1) limited availability of labelled training data, due to the expertise and excessive time required to create annotations; 2) class imbalance, particularly where the class of interest (e.g., tumour) is infrequently represented in the dataset; 3) complex and ambiguous features which are difficult to identify from Magnetic Resonance Imaging (MRI) data with low signal to noise ratio, leading to variation in annotations from different graders for the same image; and 4) considerable anatomic variability between patients. As a result, quantifying the uncertainty in predicted segmentations, regardless of the chosen method, is paramount to prevent over-confident and biased diagnoses czolbe2021segmentation ; varoquaux2022machine . Generating multiple plausible segmentations also supports ensemble-based uncertainty quantification in biomedical computational fluid dynamics (CFD) models (e.g., gaidzik2021hemodynamic ; chen2020deep ). We propose an accurate and computationally feasible method for generating ensembles of plausible segmentations in medical applications plagued by the aforementioned issues.
@@
Significant research effort in the field of deep learning based medical image segmentation has focused on improving the accuracy of convolutional neural networks (CNNs). By now, the U-Net Ronneberger2015 and Attention U-Net oktay2018attention have emerged as empirically successful methods for image segmentation in a range of computer vision tasks not limited to medical imaginEnhancements of such methods have been developed to remain competitive when trained on small data sets, e.g., via deformation consistent, multiscale supervision Chatterjee2020 , along with improved loss functions for heavily class imbalanced data sets, such as the Focal Tversky Loss abraham2019novel . However, uncertainty quantification of such outputs remains a considerable challenge, which is exacerbated when training data are sparse. Furthermore, the logits or activations produced from most deep learning based segmentation methods provide only a pixel-wise measure of uncertainty and cannot be reliably used to generate samples from the underlying joint distribution Monteiro2020 ; krygier2021quantifying . This necessitates tailor-made methods that carefully account for the joint probability between pixels, as in, for instance, Deep Ensembles caldeira2020deeply and Stochastic Segmentation Networks Monteiro2020 .
@@
While Bayesian methods li2021deep ; abdar2021review ; kwon2020 are supported by rich theory for quantifying both aleatoric and epistemic uncertainty, it is necessary to resort to approximations of both the prior and posterior in most high dimensional medical imaging applications. In Bayesian neural networks, weights are treated as random variables whose distributional properties must also be estimated from data. This is particularly challenging in medical imaging problems where annotated data sets are limited, thus exacerbating the issue of parameter identifiability khemakhem2020variational . Arguably the most prominent, flexible and easily implementable approach is Monte Carlo dropout Gal2016 ; Mobiny2021 , which involves randomly removing elements of the input tensor hinton2012improving or removing channels (i.e., feature maps) while using spatial dropouts with convolutional layers tompson2015efficient , during both training and inference. This serves as a fast but somewhat crude approximation to the posterior. Variational autoencoder and variational inference methods, such as the Probabilistic U-Net kohl2018 , Variational Inference U-Net (VI U-Net) and Multi-Head VI U-Net Fuchs2022 ) can serve as a fast and reasonably approximation, as they not only drastically simplify computation of the posterior, but typically also capture uncertainty through a lower dimensional latent variable which is more identifiable. However, these methods rely on a cross-entropy style reconstruction term in the evidence lower bound (ELBO), which is known to be problematic for class imbalanced data sets abraham2019novel ; rezaei2020addressing ; tian2022striking , as is common in medical image segmentation.
A modified version of the Probabilistic U-net kohl2018 that allows for improved representation of inter-rater variability in annotations in class-imbalanced data has been investigated in this research, by experimenting with a range of distribution losses. The literature lacks 3D implementation of generative models, whereas a comparison between 2D and 3D models is rarely performed. Hence, this research also investigates the influence of training on 2D slices compared to 3D patchesTesting of the method is primarily focused on intracranial segmentations of blood vessels and lesions due to multiple sclerosis. The features of interest are complex in their geometry and ambiguous and there exists appreciable variability in annotations between graders. We focus on methods capable of generating ensembles of segmentations that can then be used to represent uncertainty in the model domain in hemodynamic models. These models are crucial for better understanding the development and diagnoses of cardiovascular and neurovascular diseases which count as the 1st and 2nd leading cause of death worldwide fan2018neurovascular ; campbell2019ischaemic ; chen2020deep . Although we verify our method on intracranial segmentation, it can be applied to other segmentation tasks where multiple annotations per image are available.
@@ 2.2 U-Net
The U-Net Ronneberger2015 is a CNN characterised by a contracting network comprising successive convolution, ReLU activation and max-pooling layers that generates low resolution feature maps, combined with an expanding network where max-pooling is replaced by upsampling to increase the resolution of the output. Skip connections allow high resolution features generated along the contracting network to be combined with the upsampling output, thereby improving the precision of the segmentation. We summarise the U-Net architecture using the following simplified notation
@@
This approach is known to produce unrealistic and noisy segmentations singh2023improving , hence we adopt a commonly used image thresholding strategy suitable for binary segmentation whereby
@@
where $\tau$ is decision threshold. This threshold is determined using Otsu’s method otsu1979threshold which gives the optimal value of $\tau$ that minimises intra-class intensity variance. Note that $\tau=0.5$ corresponds to (6).
@@ 2.3 Probabilistic U-Net
The U-Net in principle provides probabilistic outputs through independent pixel-wise categorical distributions (see 3), although it provides a poor representation of the dependence between pixels. The Probabilistic U-Net kohl2018 has emerged as an empirically competitive approach to quantifying aleatoric uncertainty by making use of ideas from conditional variational auto-encoders (cVAEs) and $\beta$-VAEs Higgins2017 ; Kingma2019 . This is done by incorporating a random latent factor $z\in\mathbb{R}^{d}$, that represents a low dimensional embedding of the image-segmentation pair, which is then used to efficiently generate samples of plausible segmentations $s$. The distribution of $z$ is captured through the intractable posterior $p_{\theta}(z|{\bf x,y})$, whose form is learned from the data. In cVAEs, the goal is to determine $\theta,\phi$ such that a more tractable form $q_{\phi}(z|{\bf x,y})$ best approximates the intractable posterior $p_{\theta}(z|{\bf x},{\bf y})$ and $p_{\theta}(x|{\bf y})$ approximates the true conditional data distribution $p(x|{\bf y})$. As is common in many applications of variational inference, the posterior $p_{\theta}(z|{\bf x,y})$ is assumed to be well approximated by a Gaussian distribution,
@@ 2.4 Multi-head Variational Inference U-Net (MH VI U-Net)
The MH-VI U-Net Fuchs2022 is a variational inference based extension of the multi-headed U-Netwhich aims to improve out-of-distribution robustness. The multi-head approach consists of a base model (e.g., a U-Net) and an ensemble of “heads”, typically a series of convolutions in the final layers of the U-Net. The MH-VI U-Net aims to improve on the classical MH U-Net by replacing regular convolutions with Bayesian convolutions in the heads. The classical MH U-Net allows only for maximum likelihood estimation of the weights at each head.

### v2

@@ Abstract
In the domain of medical imaging, many supervised learning based methods for segmentation face several challenges such as high variability in annotations from multiple experts, paucity of labelled data and class imbalanced datasets. These issues may result in segmentations that lack the requisite precision for clinical analysis and can be misleadingly overconfident without associated uncertainty quantification. This work proposes the PULASki method as a computationally efficient generative tool for biomedical image segmentation that accurately captures variability in expert annotations, even in small datasets. This approach makes use of an improved loss function based on statistical distances in a conditional variational autoencoder structure (Probabilistic UNet), which improves learning of the conditional decoder compared to the standard cross-entropy particularly in class imbalanced problems. The proposed method was analysed for two structurally different segmentation tasks (intracranial vessel and multiple sclerosis (MS) lesion) and compare our results to four well-established baselines in terms of quantitative metrics and qualitative output. These experiments involve class-imbalanced datasets characterised by challenging features, including suboptimal signal-to-noise ratios and high ambiguity. Empirical results demonstrate the PULASKi method outperforms all baselines at the 5% significance level. Our experiments are also of the first to present a comparative study of the computationally feasible segmentation of complex geometries using 3D patches and the traditional use of 2D slices. The generated segmentations are shown to be much more anatomically plausible than in the 2D case, particularly for the vessel task. Our method can also be applied to a wide range of multi-label segmentation tasks and and is useful for downstream tasks such as hemodynamic modelling (computational fluid dynamics and data assimilation), clinical decision making, and treatment planning.
@@ 1 Introduction
The recent explosion in deep learning based medical image segmentation techniques has led to a dramatic improvement in the rapid diagnosis, treatment planning and modelling of various diseases [1, 2]. Medical image segmentation has specific challenges compared to other classical computer vision tasks, which has spurred the development of several bespoke machine learning tools [3, 4, 5]. Such issues include 1) limited availability of labelled training data, due to the expertise and excessive time required to create annotations; 2) class imbalance, particularly where the class of interest (e.g., tumour) is infrequently represented in the dataset; 3) complex and ambiguous features which are difficult to identify from Magnetic Resonance Imaging (MRI) data with low signal to noise ratio, leading to variation in annotations from different graders for the same image; and 4) considerable anatomic variability between patients. As a result, quantifying uncertainty in predicted segmentations, regardless of the chosen method, is paramount to prevent over-confident and biased diagnoses [6, 7]. Generating multiple plausible segmentations also supports Monte-Carlo based uncertainty quantification in biomedical computational fluid dynamics (CFD) modelling (e.g., [8, 9]). CFD based modelling of the onset and progression of cardiovascular and neurovascular diseases such as aneurysms requires capturing uncertainty in various components of the modelling procedure, so as to not produce over-confident predictions. This work proposes an accurate and computationally feasible method for generating multiple plausible segmentations from a single image in medical applications plagued by the aforementioned issues.
@@
Significant research effort to date has focused on improving the accuracy of deep-learning based medical image segmentation, e.g. using convolutional neural networks (CNNs). By now, the U-Net [10] and Attention U-Net [11] have emerged as empirically successful methods for image segmentation in a range of computer vision tasks not limited to medical imaging. Enhancements of such methods have been developed to remain competitive when trained on small data sets, e.g., via deformation consistent, multiscale supervision [12], along with improved loss functions for heavily class imbalanced data sets, such as the Focal Tversky Loss [13]. However, uncertainty quantification of such outputs remains a considerable challenge, which is exacerbated when training data are sparse. Furthermore, the logits or activations produced from most deep learning based segmentation methods provide only a pixel-wise measure of uncertainty and cannot be reliably used to generate samples from the underlying joint distribution [14, 15]. This necessitates tailor-made methods that carefully account for the joint probability between pixels, as in, for instance, Deep Ensembles [16] and Stochastic Segmentation Networks [14].
@@
One of the most thoroughly investigated methods for uncertainty quantification in deep learning is with the Bayesian framework. In the Bayesian approach, prior knowledge of one or more components of the architecture (e.g. neural network weights) are combined with data via Bayes theorem. While Bayesian methods [17, 18, 19] are supported by rich theory for quantifying both aleatoric and epistemic uncertainty, it is necessary to resort to approximations of both the prior and posterior distributions in most high dimensional medical imaging applications. Arguably the most prominent, flexible and easily implementable approach is Monte Carlo dropout [20, 21], which involves randomly removing elements of the input tensor [22] or removing channels (i.e., feature maps) while using spatial dropouts with convolutional layers [23], during both training and inference. This serves as a fast but somewhat crude approximation to the posterior. In Bayesian neural networks, weights are treated as random variables whose distributional properties must also be estimated from data. This is particularly challenging in medical imaging problems where annotated data sets are limited, thus exacerbating the issue of parameter identifiability [24]. More recently, Variational inference has emerged as a competitive tool for fast approximate Bayesian computations. Variational autoencoder and variational inference methods, such as the Probabilistic U-Net [25] which is a major focus of this paper, Variational Inference U-Net (VI U-Net) and Multi-Head VI U-Net [26]) employ a simpler parametric distribution to approximate the posterior (e.g. Normal) to drastically simplify computations and typically also capture uncertainty through a lower dimensional latent variable which is more identifiable. However, these methods rely on a cross-entropy style reconstruction term in the evidence lower bound (ELBO), which is known to be problematic for class imbalanced data sets [13, 27, 28], as is common in medical image segmentation. This paper proposes a modification of this loss function to improve generation of plausible segmentations for class imbalanced datasets with considerable inter-rater variability.
Another important consideration in generating segmentations, particularly in medical applications where images contain complex intricate details, is whether to use 2D slices or 3D volumes. When generating 3D volumes, it is common to use smaller 3D patches when a full 3D model is computationally infeasible due to resource limitations. While 3D patches typically provide superior encoding of spatial relationships, they come with increased computational complexity. Despite the clear benefits of 3D models in capturing the spatial structures, there is a notable gap in the literature regarding the use of 3D implementations for generative models, and direct comparisons between 2D and 3D models remain rare. A second goal of this paper is to address this gap by investigating how training on 2D slices compares to training on 3D patches, offering valuable insights into which approach may be more effective, particularly in medical imaging applications.
1.1 Contributions
This paper introduces the PULASki method, a computationally efficient generative tool for biomedical image segmentation that effectively captures variability in expert annotations, even within small datasets. Our method relies on an easily implementable and interpretable modification of popular probabilistic U-nets [25] primarily via improved loss functions. These loss functions are based on statistical distances, thereby enhancing the learning process of well-established probabilistic U-Nets by incorporating the complete set of annotations from multiple raters simultaneously, rather than learning from a single random annotation at a time (see Sec. 3). Incorporating statistical distance measures into loss functions can enhance model performance on imbalanced datasets by enabling a more nuanced comparison between predicted and actual data distributions. We consider a range of distances that improve the model’s ability to accurately represent minority classes (and local geometric features), while also allowing the model to focus on the underlying data distribution, leading to better generalisation and performance in scenarios characterised by class imbalance. The advantage of our method is its potential to be easily taken up by practitioners given its structural similarity to the U-net and its popularity for image segmentation. A major focus here is in testing our method on datasets characterising real-world challenges such as: small sample sizes, class-imbalance, high ambiguity and memory requirements for 3D volumes. We evaluate PULASki on intracranial vessel and multiple sclerosis lesion segmentation tasks, demonstrating superior performance compared to established baselines. The features of interest are complex in their geometry and ambiguous and there exists appreciable variability in annotations between graders. Furthermore, the paper presents a comparative study of 2D and 3D segmentation models. Although the experiments focused on two specific segmentation tasks, the method is applicable to other segmentation tasks characterised by high inter-rater disagreement. The versatility of the PULASki method has the potential to be extended to various multi-label segmentation tasks, supporting applications in clinical decision-making, treatment planning, and haemodynamic modelling to better understand cardiovascular and neurovascular diseases which are leading causes of death worldwide [29, 30, 9].
@@ 2.2 U-Net
The U-Net [10] is a CNN characterised by a contracting network comprising successive convolution, ReLU activation and max-pooling layers that generates low resolution feature maps, combined with an expanding network where max-pooling is replaced by upsampling to increase the resolution of the output. Skip connections allow high resolution features generated along the contracting network to be combined with the upsampling output, thereby improving the precision of the segmentation. We summarise the U-Net architecture using the following simplified notation
@@
This approach is known to produce unrealistic and noisy segmentations [31], hence we adopt a commonly used image thresholding strategy suitable for binary segmentation whereby
@@
where $\tau$ is decision threshold. This threshold is determined using Otsu’s method [32] which gives the optimal value of $\tau$ that minimises intra-class intensity variance. Note that $\tau=0.5$ corresponds to (6).
@@ 2.3 Probabilistic U-Net
The U-Net in principle provides probabilistic outputs through independent pixel-wise categorical distributions (see 3), although it provides a poor representation of the dependence between pixels. The Probabilistic U-Net [25] has emerged as an empirically competitive approach to quantifying aleatoric uncertainty by making use of ideas from conditional variational auto-encoders (cVAEs) and $\beta$-VAEs [33, 34]. This is done by incorporating a random latent factor $z\in\mathbb{R}^{d}$, that represents a low dimensional embedding of the image-segmentation pair, which is then used to efficiently generate samples of plausible segmentations $s$. The distribution of $z$ is captured through the intractable posterior $p_{\theta}(z|{\bf x,y})$, whose form is learned from the data. In cVAEs, the goal is to determine $\theta,\phi$ such that a more tractable form $q_{\phi}(z|{\bf x,y})$ best approximates the intractable posterior $p_{\theta}(z|{\bf x},{\bf y})$ and $p_{\theta}(x|{\bf y})$ approximates the true conditional data distribution $p(x|{\bf y})$. As is common in many applications of variational inference, the posterior $p_{\theta}(z|{\bf x,y})$ is assumed to be well approximated by a Gaussian distribution,
@@
2.4 Diversified and Personalised Multi-rater Segmentation (D-Persona)
D-Persona [35] was proposed to address the challenges of multi-rater medical image segmentation, where annotation ambiguity arises due to data uncertainties and diverse observer preferences. Unlike traditional methods, D-Persona aims to generate both diversified and personalised segmentation results by employing a novel two-stage framework.
Stage I: Diversified Segmentation
The first stage learns a common latent space to capture the variability across multiple expert annotations. Using a Probabilistic U-Net as the base model, this stage introduces a bound-constrained loss to enhance segmentation diversity.
The latent space is defined by prior and posterior distributions:
$\mu_{\text{prior}},\sigma_{\text{prior}}$ $=F_{\text{prior}}(X),$ (17)
$\mu_{\text{post}},\sigma_{\text{post}}$ $=F_{\text{post}}(X,\mathcal{A}_{\text{set}}),$
where $F_{\text{prior}}$ and $F_{\text{post}}$ are neural encoders, $X$ represents the input, and $\mathcal{A}_{\text{set}}$ denotes the set of expert annotations. The Kullback-Leibler divergence aligns these distributions:
$L_{\text{KL}}=\text{KL}(D_{\text{prior}}(X)\|D_{\text{post}}(X,\mathcal{A}_{\text{set}})).$ (18)
To encourage diversity, intersections and unions of annotations ($\mathcal{A}_{\text{bound}}=\{\mathcal{A}_{\text{inter}},\mathcal{A}_{\text{union}}\}$) are computed, and a bound loss is applied:
$L_{\text{bound}}=\text{DSC}(P_{\text{prior, inter}},\mathcal{A}_{\text{inter}})+\text{DSC}(P_{\text{prior, union}},\mathcal{A}_{\text{union}}),$ (19)
where $P_{\text{prior, inter}}$ and $P_{\text{prior, union}}$ are predictions derived from the prior distribution. The overall loss for Stage I is:
$L_{\text{stage 1}}=L_{\text{KL}}+\alpha L_{\text{seg}}+\beta L_{\text{bound}},$ (20)
where $L_{\text{seg}}$ is a segmentation loss based on randomly sampled latent codes.
Stage II: Personalised Segmentation
In the second stage, D-Persona generates personalised segmentations by learning expert-specific prompts from the common latent space. For each expert $i$, a projection head $F_{\text{proj},i}$ maps the U-Net features into expert-specific latent codes:
$z_{\text{exp},i}=\text{Pooling}[F_{\text{proj},i}(F_{\text{backbone}}(X))].$ (21)
To ensure these latent codes remain within the learned latent space, a cross-attention mechanism aligns $z_{\text{exp},i}$ with a prior bank:
$\tilde{z}_{\text{exp},i}=\text{Softmax}(z_{\text{exp},i}^{\top}\cdot z_{\text{prior bank}})\cdot z_{\text{prior bank}}^{\top}.$ (22)
The final segmentation for expert $i$ is obtained by:
$P_{\text{exp},i}=F_{\text{head}}(\tilde{z}_{\text{exp},i},F_{\text{backbone}}(X)),$ (23)
and supervised using:
$L_{\text{corr, seg}}=\sum_{i=1}^{n}\text{DSC}(P_{\text{exp},i},A_{i}),$ (24)
where $A_{i}$ is the annotation corresponding to expert $i$.
2.5 Collectively Intelligent Medical Diffusion (CIMD)
Collectively Intelligent Medical Diffusion (CIMD) [36] was proposed to address the ambiguity inherent in medical image segmentation by utilising the stochastic nature of diffusion models. Unlike the Probabilistic U-Net and its hierarchical variants, CIMD generates multiple plausible segmentation masks without the need for an additional network to encode prior distributions. The method directly models ambiguity during training using a novel hierarchical diffusion framework.
The diffusion process consists of two phases: a forward process that iteratively adds Gaussian noise to the segmentation mask $x_{0}$, and a reverse process that denoises the perturbed mask to recover clean annotations:
$q(x_{t}|x_{t-1})$ $=\mathcal{N}(x_{t};\sqrt{\alpha_{t}}x_{t-1},(1-\alpha_{t})I),$ (25)
$p_{\theta}(x_{t-1}|x_{t})$ $=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$
where $\alpha_{t}$ is a noise scheduling parameter, and $\mu_{\theta}$ and $\Sigma_{\theta}$ are learned by the model.
To capture the ambiguity of medical segmentations, two neural networks are used:
  1. 1.
    The Ambiguity Modelling Network (AMN), which models the distribution of ground truth annotations as:
    $Q(z|b,x_{b})=\mathcal{N}(\mu(b,x_{b};\nu),\sigma(b,x_{b};\nu)),$ (26)
    where $\nu$ represents trainable parameters.
  2. 2.
    The Ambiguity Controlling Network (ACN), which models the ambiguity of predictions during the reverse process as:
    $P(z|b,\hat{x}_{b})=\mathcal{N}(\mu(b,\hat{x}_{b};\omega),\sigma(b,\hat{x}_{b};\omega)),$ (27)
    where $\omega$ denotes trainable parameters.
The loss function combines components to optimise the reverse process and align the ambiguity distributions of AMN and ACN:
$L_{\text{total}}=L_{\text{simple}}+\lambda L_{\text{vlb}}+\beta D_{\text{KL}}(Q(z|b,x_{b})\|P(z|b,\hat{x}_{b})),$ (28)
where $L_{\text{simple}}$ minimises the noise prediction error, $L_{\text{vlb}}$ is a variational lower bound, and the KL divergence term $D_{\text{KL}}$ penalises the mismatch between AMN and ACN distributions.
Unlike traditional models, CIMD introduces stochasticity at each hierarchical level, enabling the generation of diverse and accurate segmentation masks. These masks align with ground truth distributions and inherently capture inter-expert variability without requiring prior networks. The sampling process during inference iteratively refines segmentation:
$x_{t-1}\leftarrow\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\gamma_{t}}}f_{\theta}(x_{t},t)\right)+\gamma_{t}z,$ (29)
where $z\sim\mathcal{N}(0,I)$ and $t\in[1,T]$.
This framework significantly improves segmentation diversity and accuracy, particularly in medical contexts requiring collective intelligence for diagnostics.
@@ 2.6 Multi-head Variational Inference U-Net (MH VI U-Net)
The MH-VI U-Net [26] is a variational inference based extension of the multi-headed U-Netwhich aims to improve out-of-distribution robustness. The multi-head approach consists of a base model (e.g., a U-Net) and an ensemble of “heads”, typically a series of convolutions in the final layers of the U-Net. The MH-VI U-Net aims to improve on the classical MH U-Net by replacing regular convolutions with Bayesian convolutions in the heads. The classical MH U-Net allows only for maximum likelihood estimation of the weights at each head.